how do you solve this -5^2 compared to(-5)^2
if there are no parenthesis around the number and the negetive, does the negetive sign get squared?

Answer by stanbon(75887)

-5^2 compared to(-5)^2
----------------
-5^2 = -5*5 = -25
-----------
(-5)^2 = (-5)(-5) = 25
=========================
Cheers,
Stan H.